The U.S. Census Bureau computed the proportion of U.S residents who live in each of the four geographic regions in 2010. Then a simple random sample was drawn of 1000 people living in the U.S. in 2018. The following table presents the results:

Region 2010 Proportion 2018 Frequency
Northeast 0.179 173
Midwest 0.217 205
South 0.371 384
West 0.233 238

Determine whether the proportions of people living in the U.S. in the various regions has changed between 2010 and 2018. Use a 0.05 level of significance.

Prompt 1: Null Hypothesis

Prompt 1 Answer

Prompt 2: Number of degrees of freedom?

Prompt 2 Answer

Prompt 3: P-value

Prompt 3 Answer

Prompt 4Has the proportion of people living in the different regions of the U.S. changed from 2010 to 2018?

prompt 4: answer
Answer #1

H0: The proportion of people living in different region in 2010 and 2018 is same.

H0: the proportion of people living in different region has changed between 2010 and 2018.

Categories Observed Expected Northeast 173 1000*0.179=179 Midwest 205 1000*0.217=217 South 384 1000*0.371=371 West 238 1000*0(1) Null and Alternative Hypotheses The following null and alternative hypotheses need to be tested: H. : P1 = 0.179, P2 = 0.(3) Test Statistics The Chi-Squared statistic is computed as follows: n (Oi x2 = Σ E;)? Ei - 0.201 + 0.664 +0.456 + 0.107 = 1(5) Conclusion It is concluded that the null hypothesis Ho is not rejected. Therefore, there is NOT enough evidence to claimthus it may concluded that the proportion of people living in different region has not changed over 2010 to 2018.

I have given my best to solve your problem. Please like the answer if you are satisfied with it. ?